A (B+25.0) g mass is hung on a spring. As a result the spring stretches (8.50+A) cm. If the object is then pulled an additional 3.00 cm downward and released, what is the period of the resulting oscillation? Give your answer in seconds with 3 significant figures.

A=9, B=081
Answer #1

even (BF 259) (81)cm, en A-9; B=081 At equilibrium point, Fnet = kx-mg=0 - kn=ng m= 081+ 259 = 1069 = alobky x 8.5+ 9 = 17.5